OpenStudy (anonymous):

Find any relative extrema of the function. (Round your answers to three decimal places.) f(x) = arcsec(x) − 3x relative maximum (x, y) = ( ) relative minimum (x, y) = ( )

OpenStudy (anonymous):

can someone please help me..

OpenStudy (anonymous):

\[f(x)=\sec^{-1}(x)-3x\] \[f'(x)=\frac{1}{x\sqrt{x^2-1}}-3\]

OpenStudy (anonymous):

set this beast equal to zero and solve

OpenStudy (anonymous):

cana u break it down for me till the simple algebra so i can solve it from there.please

OpenStudy (anonymous):

ok if i can. is the derivative ok?

OpenStudy (anonymous):

\[f'(x)=\frac{1}{x\sqrt{x^2-1}}-3=0\] \[\frac{1}{x\sqrt{x^2-1}}=3\] \[3x\sqrt{x^2-1}=1\] \[9x^2(x^2-1)=1\] \[9x^4-9x^2-1=0\]

OpenStudy (anonymous):

now you have a quadratic equation is \[x^2\] use the quadratic formula and solve for z in \[9z^2-9z-1=0\]

OpenStudy (anonymous):

or just cheat. may be easier http://www.wolframalpha.com/input/?i=y%3Darcsec%283x%29-3x

OpenStudy (anonymous):

satelitte i clciked on the link but what would be my relative maximum and relative minimum...could you point that out to me

OpenStudy (anonymous):

it is at the very bottom of the link

OpenStudy (anonymous):

max at x = stuff min at x = stuff

OpenStudy (anonymous):

but doesnt it have to be a point like (x,y) so that solves for x what about y.

OpenStudy (anonymous):

sorry to be such a pain but could you help me get two a solid point for the min and maximum

OpenStudy (anonymous):

oh wait i think i got it

OpenStudy (anonymous):

ugh you plugged in 3x instead of just x into wolfram so i got it wrong its oki though

OpenStudy (anonymous):

could u help me with something esle

OpenStudy (anonymous):

but is said 3x in your problem yes?

OpenStudy (anonymous):

i just used your original equation

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=f%28x%29%3Darcsec%28x%29-3x i typed in just as you wrote it